3n(7n-20)=96
i have gotten to
21n squared -60=96
then im stuck :(
please expain how to work this or at least show the steps.

Answer by edjones(8007)

3n(7n-20)=96
21n^2-60n=96
21n^2-60n-96=0
21*-96=-2016
what two factors of -2016 have a sum of -60? -84 and 24.
21n^2+24x-84x-96=0
3n(7n+8)-12(7n+8)=0 Factoring by grouping.
(3n-12)(7n+8)=0
.
3n=12
n=4
.
7n=-8
n=-8/7
.
Ed